# A new characterization of (pre)liminary C*-algebras

### v3

@@ 7.1. The Pedersen weight
(1) follows from [4, Corollary 2.5].
@@
for all $n\in\mathbb{N}$. Let $h\left(b,A\right)=\overline{b^{\ast}Ab}$ be the hereditary C*-subalgebra of $A$ generated by $b$. Recall the correspondence between irreducible representations of $h\left(b,A\right)$ and irreducible representations of $A$ that do not vanish on $h\left(b,A\right)$ [48, Section 4.1]. By this correspondence, without loss of generality, one can replace $A$ with $h\left(b,A\right)$. After doing so, one obtains that $\mathrm{\dim}\left(\rho_{n}\right)=d$ for every $n\in\mathbb{N}$. In turn, this implies by [4, Corollary 2.5] that $C=\left\{\mathrm{\mathrm{Ker}}\left(\pi_{1}\right),\ldots,\mathrm{\mathrm{Ker}}\left(\pi_{\ell}\right)\right\}$ is a finite set with at most $n$ elements. Furthermore,
@@
7.2. Fell’s condition of order $\infty$
Let $A$ be a separable C*-algebra with primitive spectrum $X$, and $\pi_{0}$ be an irreducible representation of $A$. Let us say that a local section for $A$ is a pair $\left(b,U\right)$ where $b\in A$, $U\subseteq X$ is open, and $\rho\left(b\right)\neq 0$ for every $\mathrm{\mathrm{Ker}}\left(\rho\right)\in U$. This is a local section around $\pi_{0}$ if $\mathrm{\mathrm{Ker}}(\pi_{0})\in U$. Recall the definition of the Pedersen weigh $w_{b}$ associated with an element $b$ of $A$ as in Definition 7.1.
Definition 7.4.
Let $A$ be a separable C*-algebra, and $\pi_{0}$ an irreducible representation of $A$. Then $\pi_{0}$ satisfies Fell’s condition of order $\infty$ if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that
$w_{b}|_{U}:U\rightarrow\overline{\mathbb{N}}\text{, }\mathrm{\mathrm{Ker}}\left(\rho\right)\mapsto\mathrm{\mathrm{rank}}\left(\rho\left(b\right)\right)$
is finite.
It turns out that such a condition is automatically satisfied whenever $\pi_{0}$ has a neighborhood of CCR representations. This is essentially established in the proof of [24, Lemma 2.5].
@@ 7.2. Fell’s condition of order $\infty$
Let $A$ be a separable C*-algebra with primitive spectrum $X$, and $\pi_{0}$ an irreducible representation of $A$. Suppose that $\mathrm{\mathrm{Ker}}(\pi_{0})$ has a neighborhood $W$ in $X$ such that every irreducible representation $\rho$ of $A$ with $\mathrm{\mathrm{Ker}}\left(\rho\right)\in W$ is CCR. Then there exists a positive element $z\in A$ and a neighborhood $U$ of $\mathrm{\mathrm{Ker}}\left(\pi_{0}\right)$ contained in $W$ such that:
@@
    $\pi_{0}\left(z\right)$ is a rank-one projection, and
@@
In particular, $\pi_{0}$ satisfies Fell’s condition of order $\infty$.
@@
Let $P\in K\left(H_{\pi_{0}}\right)$ be a rank-one projection. Pick a positive element $z\in A$ such that $\pi_{0}\left(z\right)=P$. Fix $n\geq 2$ and set $\varepsilon:=1/n$. Let $F$ be a continuous non-negative function on $\mathbb{R}$ which satisfies
@@
$\pi_{0}\left(F\left(z\right)\right)=F(\pi_{0}\left(z\right))=F\left(P\right)=P\text{.}$
Considering the open condition
$\left\|F\left(z\right)+J\right\|>1-\varepsilon$
we see that there exists a neighborhood $U$ of $\mathrm{\mathrm{Ker}}\left(\pi_{0}\right)$ contained in $W$ such that
$\left\|F\left(z\right)+J\right\|>1-\varepsilon$
@@
$F\left(z\right)+J\neq 0$
for all $J\in U$. By assumption, every irreducible representation $\rho$ with $\mathrm{\mathrm{Ker}}\left(\rho\right)\in W$ is CCR. Thus, for $J\in U$ we have that $F\left(z\right)+J$ is a compact positive operator with all nonzero eigenvalues in $1-\varepsilon,1+\varepsilon]$. This implies that $F\left(z\right)+J$ has finite rank.
This shows that $\left(z,U\right)$ is a local section around $\pi_{0}$ such that $w_{z}|_{U}$ is finite, and it is a weight by Proposition 7.5. ∎
Recall that an irreducible representation $\pi_{0}$ of a separable C*-algebra satisfies Fell’s condition of order $1$ if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that the weight $w_{b}|_{U}$ is finite of rank at most $1$. More generally, $\pi_{0}$ satisfies Fell’s condition of order $k<\omega$ as in [21, Section 3.2] if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that the weight $w_{b}|_{U}$ is finite of rank at most $k$.
@@ 7.4. Pedersen order of positive elements
Recall that an element $x$ of a separable C*-algebra $A$ is abelian if and only if the hereditary C*-algebra $\overline{x^{\ast}Ax}$ is commutative. This is equivalent to the assertion that $\mathrm{\mathrm{rank}}\left(\pi\left(x\right)\right)\leq 1$ for every irreducible representation $\pi$ of $A$, i.e., the Pedersen weight $w_{x}$ is finite of rank at most $1$. More generally, $x$ has global rank at most $k<\omega$ as in [21, Section 3.2] if $w_{x}$ is finite of rank at most $k$. We consider the following generalization of these notions:
@@
If $B$ is a hereditary C*-subalgebra of $A$, then every irreducible representation of $B$ is unitarily equivalent, for some irreducible representation $\pi$ of $A$ whose kernel does not contain $B$ [48, Proposition 4.1.9], to the restriction of $\pi|_{B}$ to its essential subspace [7, II.6.1.5]. Therefore, if $b\in B$ then its Pedersen rank is the same when computed in $B$ or in $A$.
@@ 7.6. Type $I_{\alpha}$ C*-algebras and $\alpha$-subhomogeneous C*-algebras
The following definition subsumes the notion of type $I_{n}$ C*-algebra for $n<\omega$ from [21, Section 3.2], which in turn has the notion of type $I_{0}$ C*-algebra as particular case. It also subsumes the notion of $n$-subhomogeneous C*-algebra, which is recovered as a particular case when $\alpha=n<\omega$.
@@ 8. Fell and Pedersen
In this section, we characterize type $I_{\alpha}$ C*-algebras in terms of the Fell rank of elements, generalizing [21, Proposition 17].
@@ 8.1. Fell rank and Pedersen rank
As a preliminary observation, we describe the Fell rank of an irreducible representation in terms of elements of a given Pedersen rank. In the case when $\alpha=n<\omega$ we recover [21, Lemma 16]. Indeed, we follow a similar argument.
@@ 8.3. Pedersen order and hereditary subalgebras
Let $A$ be a separable C*-algebra, and $x\in A$. The hereditary C*-subalgebra generated by $x$ is $\overline{x^{\ast}Ax}$ [7, Proposition II.3.4.2]. The following characterization of the Pedersen rank of an element generalizes [7, IV.1.1.7], which can be seen as the particular instance in the case $\alpha=1$, and [21, Lemma 16], which can be seen as the particular case when $\alpha=n<\omega$.
@@
from [48, Proposition 4.1.8 and Proposition 4.1.9]. This is obtained by assigning to an irreducible representations $\pi$ of $B$ an irreducible representations $\rho_{\pi}$ of $A$ such that $\pi$ is equivalent to the restriction of $\rho_{\pi}|_{B}$ to its essential subspace [48, Proposition 4.1.8]. Then we have that for every $a\in A$ and irreducible representation $\pi$ of $B$,
@@
as in [48, Proposition 4.1.8 and Proposition 4.1.9]. This obtained by assigning to an irreducible representation $\rho$ of $A$ that does not vanish on $B$, the restriction of $\rho|_{B}$ on its essential subspace; see also [7, Proposition II.6.1.9]. Then we have that for $\mathrm{\mathrm{Ker}}\left(\rho\right)\in X\setminus\mathrm{hull}\left(B\right)$,
@@ 9.2. Liminary C*-algebras
is open in the Jacobson topology of $X$. The union of $X_{\alpha}$ for $\alpha<\omega_{1}$ is equal to $X$. By [34, Theorem 6.9], it follows that there exists $\alpha<\omega_{1}$ such that $X=X_{\alpha}$, and hence $r_{F}\left(A\right)\leq\alpha$.
@@ 10. Arbitrarily high rank
In this section we present, for any countable ordinal, examples of separable C*-algebra of Fell rank and Pedersen rank $\alpha$. For each limit ordinal $\lambda$ we fix an increasing sequence $\left(\lambda_{n}\right)$ of successor ordinals converging to $\lambda$. We also set $\alpha_{n}:=\alpha-1$ if $\alpha$ is a successor.
@@ 10.1. Bi-unitization
In a similar fashion as in the case of the unitization, one can define a noncommutative analogue of the two-point compactification $X^{\dagger\dagger}$ of a countably compactly based $T_{0}$ space $X$ as considered in [37]. If $A$ is a C*-algebra, then we let $A^{\dagger\dagger}$ be the algebra $M_{2}\left(A\right)\oplus D_{2}\left(\mathbb{C}\right)$ where $D_{2}\left(\mathbb{C}\right)\subseteq M_{2}\left(\mathbb{C}\right)$ is the algebra of diagonal matrices, with coordinate-wise addition and scalar multiplication. The multiplication is induced by the $D_{2}\left(\mathbb{C}\right)$-bimodule structure on $M_{2}\left(A\right)$. Likewise, the C*-norm is the operator norm induced by letting $A^{\dagger\dagger}$ act as multiplication operators on $M_{2}\left(A\right)$. Equivalently, fixing a faithful nondegenerate representation $\subseteq B\left(H\right), AA^{\dagger\dagger}$ is the C*-subalgebra of $B\left(H\right)\otimes M_{2}$ generated by $A\otimes 1$ and $1\otimes D_{2}\left(\mathbb{C}\right)$. We let $\infty_{A}$ and $\infty_{A}^{\prime}$ be the distinguished characters of $A^{\dagger\dagger}$ corresponding to the points at infinity of $X^{\dagger\dagger}$, which we call characters at infinity. The projection $p_{A}\in A^{\dagger\dagger}$ correspond to a matrix unit in $D_{2}\left(\mathbb{C}\right)$. This turns $A^{\dagger\dagger}$ into a bipointed C*-algebra.
[…]

### v4

@@ 7.1. The Pedersen weight
(1) follows from [5, Corollary 2.5].
@@
for all $n\in\mathbb{N}$. Let $h\left(b,A\right)=\overline{b^{\ast}Ab}$ be the hereditary C*-subalgebra of $A$ generated by $b$. Recall the correspondence between irreducible representations of $h\left(b,A\right)$ and irreducible representations of $A$ that do not vanish on $h\left(b,A\right)$ [60, Section 4.1]. By this correspondence, without loss of generality, one can replace $A$ with $h\left(b,A\right)$. After doing so, one obtains that $\mathrm{\dim}\left(\rho_{n}\right)=d$ for every $n\in\mathbb{N}$. In turn, this implies by [5, Corollary 2.5] that $C=\left\{\mathrm{\mathrm{Ker}}\left(\pi_{1}\right),\ldots,\mathrm{\mathrm{Ker}}\left(\pi_{\ell}\right)\right\}$ is a finite set with at most $n$ elements. Furthermore,
@@
7.2. Local sections
Let $A$ be a separable C*-algebra with primitive spectrum $X$. For $t\in X$, we let $\pi_{t}$ be the corresponding irreducible representation. Let us say that a local section for $A$ is a pair $\left(b,U\right)$ where $b\in A$, $U\subseteq X$ is open, and $\pi_{s}\left(b\right)\neq 0$ for every $s\in U$. This is a local section around $t$ if $t\in U$.
Recall that $\pi_{t}$ satisfies Fell’s condition of order $1$ if there exists a local section $\left(b,U\right)$ around $t$ such that the weight $w_{b}|_{U}$ is finite of rank at most $1$. More generally, $\pi_{t}$ satisfies Fell’s condition of order $k<\omega$ as in [24, Section 3.2] if there exists a local section $\left(b,U\right)$ around $t$ such that the weight $w_{b}|_{U}$ is finite of rank at most $k$.
The following result is essentially established in the proof of [27, Lemma 2.5].
@@ 7.2. Local sections
Let $A$ be a separable C*-algebra with primitive spectrum $X$, and $t\in X$. Suppose that $t$ has a neighborhood $W$ in $X$ such that every $s\in W$, $\pi_{s}$ is CCR. Then there exists a positive element $z\in A$ and a neighborhood $U$ of $t$ contained in $W$ such that:
@@
    $\pi_{t}\left(z\right)$ is a rank-one projection, and
@@
Let $P\in K\left(H_{\pi_{0}}\right)$ be a rank-one projection. Pick a positive element $z\in A$ such that $\pi_{t}\left(z\right)=P$. Fix $n\geq 2$ and set $\varepsilon:=1/n$. Let $F$ be a continuous non-negative function on $\mathbb{R}$ which satisfies
@@
$\pi_{t}\left(F\left(z\right)\right)=F(\pi_{t}\left(z\right))=F\left(P\right)=P\text{.}$
By lower semi-continuity of
$t\mapsto\left\|\pi_{t}\left(F\left(z\right)\right)\right\|\text{,}$
there exists a neighborhood $U$ of $\mathrm{\mathrm{Ker}}\left(\pi_{t}\right)$ contained in $W$ such that
$\pi_{s}\left(F\left(z\right)\right)>1-\varepsilon$
@@
$\pi_{s}\left(F\left(z\right)\right)\neq 0$
for all $s\in U$. By assumption, for every $s\in W$, $\pi_{s}$ is CCR. Thus, for $s\in U$ we have that $\pi_{s}\left(F\left(z\right)\right)$ is a compact positive operator with all nonzero eigenvalues in $1-\varepsilon,1+\varepsilon]$. This implies that $\pi_{s}\left(F\left(z\right)\right)$ has finite rank.
This shows that $\left(z,U\right)$ is a local section around $t$ such that $w_{z}|_{U}$ is finite. ∎
@@ 7.4. Pedersen order of positive elements
Recall that an element $x$ of a separable C*-algebra $A$ is abelian if and only if the hereditary C*-algebra $\overline{x^{\ast}Ax}$ is commutative. This is equivalent to the assertion that $\mathrm{\mathrm{rank}}\left(\pi\left(x\right)\right)\leq 1$ for every irreducible representation $\pi$ of $A$, i.e., the Pedersen weight $w_{x}$ is finite of rank at most $1$. More generally, $x$ has global rank at most $k<\omega$ as in [24, Section 3.2] if $w_{x}$ is finite of rank at most $k$. We consider the following generalization of these notions:
@@
If $B$ is a hereditary C*-subalgebra of $A$, then every irreducible representation of $B$ is unitarily equivalent, for some irreducible representation $\pi$ of $A$ whose kernel does not contain $B$ [60, Proposition 4.1.9], to the restriction of $\pi|_{B}$ to its essential subspace [9, II.6.1.5]. Therefore, if $b\in B$ then its Pedersen rank is the same when computed in $B$ or in $A$.
@@ 7.6. Type $I_{\alpha}$ C*-algebras and $\alpha$-subhomogeneous C*-algebras
The following definition subsumes the notion of type $I_{n}$ C*-algebra for $n<\omega$ from [24, Section 3.2], which in turn has the notion of type $I_{0}$ C*-algebra as particular case. It also subsumes the notion of $n$-subhomogeneous C*-algebra, which is recovered as a particular case when $\alpha=n<\omega$.
@@ 8. Fell and Pedersen
In this section, we characterize type $I_{\alpha}$ C*-algebras in terms of the Fell rank of elements, generalizing [24, Proposition 17].
@@ 8.1. Fell rank and Pedersen rank
As a preliminary observation, we describe the Fell rank of an irreducible representation in terms of elements of a given Pedersen rank. In the case when $\alpha=n<\omega$ we recover [24, Lemma 16]. Indeed, we follow a similar argument.
@@
$\pi_{s}\left(a\right)=0\leq\pi_{s}\left(b\right)$
for any $s\in X\setminus U$, and
$\mathrm{\mathrm{rank}}\left(\pi_{s}\left(a\right)\right)\leq\mathrm{\mathrm{rank}}\left(\pi_{s}\left(b\right)\right)$
for every $s\in U$. This implies that
@@ 8.3. Pedersen order and hereditary subalgebras
Let $A$ be a separable C*-algebra, and $x\in A$. The hereditary C*-subalgebra generated by $x$ is $\overline{x^{\ast}Ax}$ [9, Proposition II.3.4.2]. The following characterization of the Pedersen rank of an element generalizes [9, IV.1.1.7], which can be seen as the particular instance in the case $\alpha=1$, and [24, Lemma 16], which can be seen as the particular case when $\alpha=n<\omega$.
@@
from [60, Proposition 4.1.8 and Proposition 4.1.9]. This is obtained by assigning to an irreducible representations $\pi$ of $B$ an irreducible representations $\rho_{\pi}$ of $A$ such that $\pi$ is equivalent to the restriction of $\rho_{\pi}|_{B}$ to its essential subspace [60, Proposition 4.1.8]. Then we have that for every $a\in A$ and irreducible representation $\pi$ of $B$,
@@
as in [60, Proposition 4.1.8 and Proposition 4.1.9]. This obtained by assigning to an irreducible representation $\rho$ of $A$ that does not vanish on $B$, the restriction of $\rho|_{B}$ on its essential subspace; see also [9, Proposition II.6.1.9]. Then we have that for $\mathrm{\mathrm{Ker}}\left(\rho\right)\in X\setminus\mathrm{hull}\left(B\right)$,
@@ 9.2. Liminary C*-algebras
is open in the Jacobson topology of $X$. The union of $X_{\alpha}$ for $\alpha<\omega_{1}$ is equal to $X$. By [43, Theorem 6.9], it follows that there exists $\alpha<\omega_{1}$ such that $X=X_{\alpha}$, and hence $r_{F}\left(A\right)\leq\alpha$.
@@
10. Trees
In this section we recall some notions about well-founded relations and trees.
10.1. Well-founded relations
A relation $R$ on a nonempty set $X$ is well-founded if every nonempty subset of $X$ has an $R$-minimal element. In this case, one can define the $R$-rank $\mathrm{rk}_{R}\left(x\right)$ of an element $x$ of $X$ by (well-founded) recursion. Thus, we have that $\mathrm{rk}_{R}\left(x\right)=0$ if and only if $x$ is $R$-minimal, and
$\mathrm{rk}_{R}\left(x\right)=\mathrm{sup}\left\{\mathrm{rk}_{R}\left(y\right)+1:yRx\right\}$
otherwise. Then
$\mathrm{rk}\left(R\right):=\mathrm{sup}\left\{\mathrm{rk}\left(x\right)+1:x\in X\right\}\text{.}$
In what follows, we will consider mainly well-founded relations that are (partial) orders, i.e., reflexive, symmetric, and transitive.
10.2. Trees
An ordered set $\left(P,\leq\right)$ is directed if for every $x,y\in P$ there exists $z\in P$ such that $x\leq z$ and $y\leq z$. A forest is an ordered set $\left(T,\leq\right)$ such that, for every $s\in T$, the set
$T_{\geq s}=\left\{t\in T:s\leq t\right\}$
of elements above $s$ is a finite linear order. A countable (rooted) tree is a countable forest that is directed. This is equivalent to the assertion that $T$ has a largest element, called the root. A tree is well-founded it is so as an ordered set, in which case its rank is defined as above. The leaves of a well-founded tree are its minimal elements, i.e., the nodes of rank 0. A countable tree is infinitely branching if every node that is not a leaf has infinitely many immediate predecessors (children). It is easily seen that any forest is a disjoint union of trees.
10.3. Topology
Given a tree, we define its canonical tree topology.
Definition 10.1.
Let $B$ a tree. A subset $O$ of $T$ is open if for every $x\in O$, $O$ contains the initial segment $T_{\leq t}$ for all but finitely many children $t$ of $x$.
Observe that, by definition, any downward-closed set is open. Given a sequence $\left(t_{n}\right)$ in $T$, we say $t=\limsup_{n}t_{n}$ if and only if $t$ is the least element of the set
$\left\{t\in T:\exists n_{0}\in\mathbb{N}\text{ such that }\forall n\geq n_{0}\text{, }t_{n}\leq t\right\}$
of elements of $T$ that majorize some tail of $\left(t_{n}\right)$. Observe that if $T_{0}$ is an subtree of $T$, then its tree topology is precisely the subspace topology inherited from $T$. The following result is easily proved by induction, and can also be seen as a consequence of the classification of scattered compact Polish spaces [43, Theorem 6.4].
Lemma 10.2.
Let $T$ be a countable well-founded infinitely-branching tree of rank $\rho+1<\omega_{1}$. Then:
  1. (1)
    $T$ is a countable compact Polish space homeomorphic to $\omega^{\rho}+1$;
  2. (2)
    a sequence $\left(t_{n}\right)$ in $T$ converges to $t$ if and only if $t=\limsup_{n}t_{n}$.
@@ 11. Arbitrarily high rank
In this section we present, for any countable ordinal, examples of separable liminary C*-algebra of Fell rank and Pedersen rank $\alpha$. In fact, the examples we present are scattered C*-algebras, and in particular approximately finite-dimensional. We will consider the case of successor ordinals, as the case of limit ordinals follows by taking direct sums. For each limit ordinal $\lambda$ we fix an increasing sequence $\left(\lambda_{n}\right)$ of successor ordinals converging to $\lambda$. We also set $\alpha_{n}:=\alpha-1$ if $\alpha$ is a successor.
@@ 11.2. The Lazar jump and Lazar C*-algebras
We notice that if $A$ is a separable preliminary C*-algebra with Hausdorff spectrum, then having Fell rank one implies having Pedersen rank one, i.e., being sum of homogeneous C*-algebras [9]. The Lazar C*-algebras show that this implication does not hold when the spectrum is not Hausdorff.
@@ 11.3. The Taylor jump and Taylor C*-algebras
11.4. C*-algebras from trees
Let $T$ be a countable well-founded infinitely-branching tree of rank $\alpha+1$. We consider $T$ as a compact Polish space with respect to the corresponding tree topology. We define a weight on $T$ by
$w_{T}:t\mapsto 2^{\mathrm{ht}\left(t\right)}$
where $\mathrm{ht}\left(t\right)$ is the height of $t$ in $T$. Then we have that $\left(T,w_{T}\right)$ is a countable, compact metrizable weighted space. In this space, the set $T_{2^{d}}^{\alpha}$ in the canonical decomposition of $\left(T,w_{T}\right)$ is the set of nodes $t\in T$ of rank $\alpha$ and height $d$. In particular, this shows that $\left(T,w_{T}\right)$ has order $\alpha+1$, and $T_{1}^{\alpha}$ only contains the root of $T$.
By recursion, we assign a unital scattered C*-algebra $A\left(T\right)$ to $T$ with $\mathrm{Prim}\left(T\right)$ isomorphic to $\left(T,w_{T}\right)$. When $\alpha=0$ define $A\left(T\right)=\mathbb{C}$. Recursively, for each child $t$ of the root of $T$, let $T_{\leq t}$ be the corresponding subtree. Define now $T$ to be the unitization of the $c_{0}$-sum of $A\left(T_{\leq t}\right)\otimes M_{2}$ where $t$ ranges among the children of the root. Then it is easily verified that $A\left(T\right)$ has the required properties.
It follows that $A\left(T\right)\otimes M_{d}\left(\mathbb{C}\right)$ is a separable unital scattered C*-algebra with spectrum homeomorphic to $\omega^{\alpha}+1$, of Fell and Pedersen rank $\omega\alpha+d$.